OpenStudy (anonymous):

The cylinder has a length of 30 centimeters and a radius is 5 centimeters. To the nearest cubic centimeter, how many cubic centimeters can one container hold? Do not round calculations until the final answer.

OpenStudy (e.mccormick):

Know the volume of a cylinder formula?

OpenStudy (anonymous):

Yes it's volume equals pi*r^2*h

OpenStudy (e.mccormick):

OK. so, what do you get when you use that?

OpenStudy (anonymous):

V=750*pi?

OpenStudy (e.mccormick):

If you think about it, a distance, or single dimension, is any single measure. An area, any area, is two measures in different directions that are multiplied, or two dimensions. A volume is three measures in three directions that are multiplied. That is the basics of any length, area, or volume. Now, take that formula: \(pi\cdot r^2\cdot h\) What you have is the area of the base, \(pi\cdot r^2\) times the height. If you had a triangular prism you would do the same basic thing. Find the area of one triangle and call that the base, then multiply by the height or length or whatever the measure of the rectangular sides of the prism are. \(30 \pi 5^2=30 \pi 25=750 \pi\) Now, they want to the nearest, so you need to multiply by pi then round.

OpenStudy (anonymous):

So would it be 2,355?

OpenStudy (e.mccormick):

Well, only if you use 3.14 for pi. If you use 3.1415 you get a different answer.

OpenStudy (anonymous):

It wants to the nearest cubic centimeter so which one should I use for pi?

OpenStudy (e.mccormick):

The more digits you use of pi, the more accurate your answer is. If you use pi in a calculator, you usually get the first 8 or 10 digits! 3.14159265359

OpenStudy (anonymous):

So what should I use

OpenStudy (anonymous):

Nevermind thank you!

OpenStudy (e.mccormick):

\(750\times 3.14 = 2355\) \(750\times 3.141 = 2355.75\) \(750\times 3.1415 = 2356.125\) \(750\times 3.1415965359 = 2356.1944901925\) As you can see, because it is 750, a few more digits of \(\pi\) make it more accurate. However, a lot more digits do not help much. For anything over 2 digits, so more than 99, adding another digit to \(\pi\) may change the answer! For over 3 digits, you can add another decimal spot to \(\pi\) again. But past about 5 digits of \(\pi\), 3.1415, is just overkill most times.

OpenStudy (anonymous):

A vertical plane passes through a cone that has a horizontal base. Which figure or figures can be produced? Select all that apply. point triangle circle line segment trapezoid rectangle

OpenStudy (e.mccormick):

OK.... hmmm.... so you have this cone:  And you are going to slice it somehow by a vertical plane. So, what do you think some of the slices might look like?

OpenStudy (e.mccormick):